The next dividend payment by CF, Inc., will be $3.00 per share. The dividends are anticipated to maintain a growth rate of 5% forever. If the stock currently sells for $20 per share, what is the required return?
Answer #1

Expected Dividend = D1 = $ 3, Current Stock Price = P0 = $ 20 and Perpetual Constant Growth Rate = g = 5 %

Let the required rate of return be R

Therefore, R = (D1/P0) + g = (3/20) + 0.05 = 0.2 or 20 %